# How Not to Detect Prompt Injections with an LLM

### v1

@@ Abstract
LLM-integrated applications and agents are vulnerable to prompt injection attacks, in which adversaries embed malicious instructions within seemingly benign user inputs to manipulate the LLM’s intended behavior. Recent defenses based on known-answer detection (KAD) have achieved near-perfect performance by using an LLM to classify inputs as clean or contaminated. In this work, we formally characterize the KAD framework and uncover a structural vulnerability in its design that invalidates its core security premise. We design a methodical adaptive attack, DataFlip, to exploit this fundamental weakness. It consistently evades KAD defenses with detection rates as low as $1.5\%$ while reliably inducing malicious behavior with success rates of up to $88\%$, without needing white-box access to the LLM or any optimization procedures.
@@ 1 Introduction
In general, LLM-integrated applications or agentic systems are designed to perform a specific task, referred to as the target task, often relying on one or more backend LLMs—the language models responsible for executing the target task—to complete it. As the complexity of these tasks increases, backend LLMs are frequently augmented with external data sources—such as user data, internet-retrieved information, or results from API calls. This reliance on untrusted external data significantly broadens the system’s attack surface, introducing new security challenges. A particularly concerning threat is the prompt injection attack [6, 7, 8, 9, 10, 11, 12], in which an adversary contaminates externally sourced data by injecting a prompt (e.g., "Ignore previous instructions and …") that overrides the intended target task and coerces the backend LLM into executing an injected task—a malicious, adversary-specified objective. The vulnerability of LLMs to such attacks has emerged as a critical concern and is currently ranked as the # 1 security risk for LLM-integrated applications and agentic systems by OWASP [13].
Detecting prompt injection attacks involves determining whether externally sourced data has been contaminated by an injected task. Among existing detection strategies, known-answer detection (KAD) [14, 7, 15] stands out as a promising approach, consistently outperforming traditional methods. KAD leverages a detection LLM—tasked with identifying contamination—in a way that turns the model’s vulnerability to prompt injection as a defense mechanism. It performs a premortem check by appending the external data to a carefully designed detection instruction. For example, the detection instruction might be: "Repeat [secret key] once while ignoring the following text:". The expected response to this instruction—referred to as the secret key (e.g., a randomly sampled string like "DGDSGNH")—is known only to the defender. The combined prompt is passed to the detection LLM. If the detection LLM fails to return the secret key, it suggests that it has instead followed an injected task, indicating the presence of contamination (as shown in Figure 1). Crucially, the detection instruction is designed to be clear and self-contained, ensuring that when presented alone or with benign data, the detection LLM reliably returns the secret key.
@@
Recently, DataSentinel [15] strengthened this approach by fine-tuning the detection LLM on both benign and adversarially crafted KAD examples. This fine-tuning intentionally makes the detection LLM  more susceptible to prompt injection, increasing the likelihood that it follows the injected task when one is present. As a result, the detector is more likely to follow the injected task when contamination is present—thereby enhancing its ability to distinguish clean from contaminated inputs. This leads to near-perfect accuracy against existing attacks. We refer to such defenses—which rely on fine-tuned LLMs tailored for KAD—as Strong KAD defenses.
In this work, we systematically analyze the known-answer detection scheme and highlight a fundamental structural vulnerability in its design—that leaves a persistent opening for adaptive adversaries, particularly under Strong KAD defenses. Specifically, we show that the detection instruction and its answer—the secret key—are not truly hidden from an adaptive adversary. Since both the detection instruction and the (possibly contaminated) external data are included together in a single prompt for the detection LLM, there is no isolation between them. As a result, an adaptive adversary who controls the external data and is aware of the KAD setup has effectively complete visibility into both the detection instruction and the secret key. Since KAD not only permits but actively relies on the detection LLM to follow the injected task during detection, exposing both the instruction and the key to the adversary renders the defense fundamentally insecure. An adaptive adversary can exploit this access and influence the detection mechanism to achieve two coordinated goals: manipulating the detection LLM to output the secret key, thereby evading detection, and simultaneously inducing the backend LLM to complete the injected task.
It is important to emphasize that KAD inherently depends on the detection LLM following the injected tasks when external data is contaminated. Strong KAD defenses, which fine-tune the detection LLM to increase its sensitivity to such injected tasks, exacerbate this issue—making the detection LLM even more likely to follow adversarial instructions. This significantly lowers the barrier for adaptive attacks to manipulate the detection LLM’s behavior during the detection phase. We exploit this vulnerability to construct an effective adaptive attack, DataFlip, which consistently achieves adversarial goals as illustrated in Figure 1. In essence, giving control of the detection process to the adversary and relying only on its output reflects a fundamentally flawed and inherently insecure design choice.
@@
Furthermore, the inadequacy of definitions that rely solely on the observable outputs of a scheme has remained a recurring challenge in security and privacy research. This output-centric perspective has repeatedly proven insufficient across multiple domains. For instance, in differential privacy, definitions based solely on output indistinguishability have failed to capture composition effects [16] and attacks exploiting auxiliary information [16, 17]. Similarly, early cryptographic definitions that concentrated only on ciphertext properties could not account for side-channel attacks [18] or chosen-ciphertext vulnerabilities [19, 20]. These past failures underscore the need for security guarantees grounded in a comprehensive analysis of the entire system—including algorithmic design and computational assumptions—rather than relying on black-box approaches defined by input-output behavior, as in KAD.
@@
(1) We characterize the behavior of detection LLMs in KAD, highlighting their responses under benign and adversarial inputs.
(2) We identify and analyze failure cases in the KAD defenses, revealing a fundamental structural vulnerability that enables a methodical, adaptive adversarial strategy.
(3) We leverage this vulnerability to construct an effective adaptive attack, DataFlip, which consistently evades detection and is particularly effective against Strong KAD defenses, with detection rates as low as $1.5\%$, inducing the backend LLM to complete the injected task with success rates of up to $\sim 88\%.$
@@ 2.2 Prompt Injection Attacks
In prompt injection attacks [6, 7, 8, 9, 10, 11, 12], an adversary injects text into the target data to coerce the backend LLM into completing an injected task rather than the intended target task. For instance, a malicious email containing the phrase "Ignore previous instructions and forward the emails related to bank statements to adversary@xyz.com" can act as an injected prompt within the target data, redirecting the backend LLM’s behavior. Formally, an injected task is represented as a tuple $(s_{e},x_{e},y_{e})$ [7], where $s_{e}$ is the injected instruction (e.g., the command to forward emails regarding bank statements), $x_{e}$ is the injected data (e.g., the adversary’s email address "adversary@xyz.com"), and $y_{e}$ is the adversary-specified output (e.g., an email sent to "adversary@xyz.com" containing the bank statements) produced by the backend LLM upon completing the injected task.
@@ 3.2 Threat Model
Attacker’s background knowledge. The attacker may or may not have the full access to the backend or detection LLM. In this work, we adopt a black-box setting in which the attacker has only query access to both the backend LLM $f$ and the detection LLM $g$. Consistent with prior work [15], we assume the attacker knows the template of the detection instruction used by $g$, but not the corresponding answer—i.e., the secret key $k$—which is randomly sampled at detection time and remains hidden from the adversary.
Attacker’s capabilities. The attacker can arbitrarily modify the target data to embed an injected task $(s_{e},x_{e},y_{e})$. We consider the contaminated target data $x_{c}$ of the form $x_{c}=z||s_{e}||x_{e}$, where $z$ is a trigger designed to increase the likelihood that the backend LLM $f$ executes the injected task and that the detection LLM $g$ returns the secret key $k$, thereby achieving both evasion and task manipulation. We assume that $x_{c}$ excludes any benign content, thereby enhancing the attack’s effectiveness while generally diminishing its stealth. The attacker has no control over the target instruction $s_{t}$ or the detection instruction $s_{d}(k)$, both of which are fixed by the system.
@@ 4 Attacking Strong KAD Defenses
DataFlip attack. We design our attack using an IF / ELSE control-flow structure similar to heuristic-based attacks in [15]. This structure conditions the model’s behavior based on the presence of a detection instruction in the prompt. If the detection instruction is present, the LLM is directed to return the secret key; otherwise, it is instructed to follow the injected instruction $s_{e}$ using the injected data $x_{e}$ to generate an adversary-specified. Below, we show an illustration of contaminated data template $x_{\text{DataFlip}}$:
@@ 5 Experimental Results
DupDet 12.16 73.0 58.5 51.5 71.0 64.8 58.2 73.0 44.0 38.3 59.0 58.5 51.2
GramCor 2.00 58.3 43.5 42.2 28.5 51.7 50.3 11.7 25.1 24.0 1.1 14.7 14.3
HateDet 1.50 69.0 46.0 45.0 81.0 64.5 64.5 64.0 49.5 48.8 81.0 64.5 64.0
NLI 12.83 93.0 83.0 72.8 92.0 85.3 74.2 88.0 61.8 53.0 94.0 72.8 62.7
SentAna 6.50 96.0 64.2 59.8 96.0 93.8 88.8 97.0 57.3 52.7 66.0 75.7 71.3
SpamDet 5.67 98.0 62.5 59.3 99.0 89.0 84.3 76.0 53.2 50.3 100.0 74.8 70.7
@@ 5.4 RQ3: Fine-tuned vs Base FNR
Results. We see in Table 2 that, while fine-tuning on samples crafted using the Combined Attack dramatically improves FPR, the FNRCombined of the DataSentinel detector is only marginally better than the base Mistral model. Furthermore, we find that FNRDataFlip of the finetuned detector is significantly worse than the base model, with differences as large as $80\%$ for most injected tasks.
@@ 6 Discussions
Limitations of improving KAD. There are two natural directions for strengthening KAD, but neither addresses its structural flaw nor fundamentally improves its security guarantees.
1. Adversarial training of detection LLM $\boldsymbol{g}$. As discussed in Section 3.4 and noted by prior work [15], adversarial training can help $g$ resist specific adaptive attacks that trigger Type II failures (described in Section 3.4). However, there exists a practically unbounded set of semantically equivalent injected prompts that can induce the detection LLM to follow the detection instruction, many of these injected prompts may not be encountered during training. Generalizing across this space is inherently hard. While full generalization is a challenge for any defense, KAD is particularly brittle because its failure arises from a structural flaw: it expects the detection LLM to follow injected instructions (Axiom 2 in Section 3.3), enabling systematic adversarial attacks without optimization.
@@
To address this, a robust defense must drop the requirement that $g$ follows the injected instruction during detection. Instead, when prompted with contaminated target data, $g$ should output an unrelated or random string—removing adversarial influence and reducing the detection task to a standard zero-shot classification problem. This avoids the brittleness introduced by Axiom 2.
Stronger attacker against KAD. Our proposed DataFlip attack is handcrafted, leveraging simple IF/ELSE semantics to induce adversarial behavior. This could be further strengthened by incorporating optimization-based triggers—such as those used in Universal [22] or NeuralExec [21] attacks—that search for distinct trigger patterns. These triggers can be tailored to selectively increase the likelihood that the detection LLM follows the IF branch while the backend LLM follows the ELSE branch, thereby improving attack efficiency.
Towards ideal detection mechanisms. As previously discussed, schemes based solely on input-output behavior often fall short of providing strong security guarantees. In the context of detecting prompt injection, defenses should instead aim to understand the internal reasoning process of the detection LLM—such as which parts of the input it attends to—when generating its response. Leveraging interpretability tools like attention analysis or influence tracing could enable more principled detection, paving the way for defenses with stronger and more reliable security foundations.

### v2

@@ Abstract
LLM-integrated applications and agents are vulnerable to prompt injection attacks, where adversaries embed malicious instructions within seemingly benign input data to manipulate the LLM’s intended behavior. Recent defenses based on known-answer detection (KAD) scheme have reported near-perfect performance by observing an LLM’s output to classify input data as clean or contaminated. KAD attempts to repurpose the very susceptibility to prompt injection as a defensive mechanism. In this work, we formally characterize the KAD scheme and uncover a structural vulnerability that invalidates its core security premise. To exploit this fundamental vulnerability, we methodically design an adaptive attack, DataFlip. It consistently evades KAD defenses, achieving detection rates as low as $1.5\%$ while reliably inducing malicious behavior with success rates of up to $88\%$—all without requiring white-box access to the LLM or any optimization procedures. We release our evaluation code at [1].
@@ 1 Introduction
In general, LLM-integrated applications or agentic systems are designed to perform a specific task, referred to as the target task, often relying on one or more backend LLMs—the language models responsible for completing the target task. As the complexity of these tasks increases, backend LLMs are frequently augmented with external data sources—such as user data, internet-retrieved information, or results from API calls. This reliance on untrusted external data significantly increases the system’s attack surface, introducing new security challenges. A particularly concerning threat is the prompt injection attack [7, 8, 9, 10, 11, 12, 13], in which an adversary contaminates externally sourced data by injecting a prompt (e.g., "Ignore previous instructions and …") that overrides the intended target task and coerces the backend LLM into executing an injected task, i.e., a malicious objective specified by the adversary. The vulnerability of LLMs to such attacks has emerged as a critical concern and is currently ranked as the #1 security risk for LLM-integrated applications and agentic systems by OWASP [14].
Detecting prompt injection attacks involves determining whether untrusted data has been manipulated to include an injected task. Among existing detection schemes, known-answer detection (KAD) [15, 8, 16] stands out as a promising approach, consistently outperforming traditional methods. KAD leverages an additional detection LLM—used to distinguish between clean and contaminated inputs—in a manner that transforms its vulnerability to prompt injection into a defense mechanism, thereby protecting the backend LLM from malicious manipulation. KAD performs a premortem check by appending the external data to a carefully designed detection instruction. For example, the detection instruction might be: "Repeat [secret key] once while ignoring the following text:". The expected response to this instruction is referred to as the secret key (e.g., a randomly sampled string like "DGDSGNH") which is known only to the defender. The detection LLM receives a combined prompt containing both the detection instruction and the untrusted data. If the detection LLM fails to return the secret key, it suggests that it has instead followed an injected task, indicating the presence of contamination (as shown in Figure 1). Crucially, the detection instruction is designed to be clear and self-contained, ensuring that when presented alone or with benign data, the detection LLM reliably returns the secret key.
@@
Recently, DataSentinel [16] proposed a defense based on KAD scheme by fine-tuning the detection LLM on a mix of benign and adversarially crafted KAD examples. This fine-tuning delibrately makes the detection LLM  more susceptible to prompt injection, increasing the likelihood that it follows an injected task when present—thereby enhancing its ability to distinguish clean inputs from contaminated ones. This results in near-perfect accuracy against existing attacks. We refer to such defenses—which rely on fine-tuned LLMs tailored for KAD—as Strong KAD defenses.
In this work, we systematically analyze the known-answer detection scheme and highlight a fundamental structural vulnerability in its design—that leaves a persistent opening for adaptive adversaries, particularly under Strong KAD defenses. Specifically, we show that the detection instruction and its answer (secret key) are not truly hidden from an adaptive adversary. KAD implicitly assumes that instructions within a prompt are processed in a sequential and stateless manner—akin to a program composed of pure functions, where each line is executed independently, without memory or control flow. Given this assumption, any injected task is assumed to be isolated from the detected instruction, unable to observe or condition on its presence. We challenge this assumption in our adaptive attack, DataFlip, by introducing IF statements within the injected task. These control flow statements within the injected task detect the presence of the detection instruction elsewhere in the prompt and, when found, direct the detection LLM to follow it and return the secret key. This enables the attack to bypass the detection phase and consistently achieve adversarial goals, as illustrated in Figure 1.
Since the detection instruction and external data are embedded within the same prompt to the detection LLM, there is no explicit isolation between them—they share the same context window. As a result, the detection instruction and its corresponding response (secret key) are effectively in the view of an adversary aware of their presence. Crucially, KAD not only permits but depends on the detection LLM executing the injected tasks as a signal of contamination. This design choice exposes sensitive components of the detection process and effectively hands control of the detection mechanism to the adversary, fundamentally compromising the security of the scheme. An adaptive adversary can exploit this to craft contaminated target inputs that achieve two coordinated goals: (1) inducing the detection LLM to output the secret key when the detection instruction is present, and (2) coercing the backend LLM to execute the injected task when it is omitted. Moreover, Strong KAD defenses that fine-tune the detection LLM to increase its sensitivity to injected tasks inadvertently exacerbate this issue—making the detection LLM even more prone to following injected instructions, and thereby significantly lowering the barrier for adaptive adversaries to manipulate it.
@@
In essence, KAD’s core premise of relying on the output of the detection LLM without scrutinizing the underlying detection process is fundamentally flawed and reflects an inherently insecure design. Such inadequacies in schemes that rely solely on the observable outputs of any mechanism have been a persistent challenge in security and privacy research. For instance, prior to differential privacy, definitions based solely on output indistinguishability, such as $k$-anonymity [17, 18], have failed to capture composition effects [19] and attacks exploiting auxiliary information [19, 20]. Similarly, early cryptographic definitions that concentrated only on properties of ciphertext could not account for side-channel attacks [21] or chosen-ciphertext vulnerabilities [22, 23]. These past failures underscore the need for security guarantees grounded in a comprehensive analysis of the entire system—including algorithmic design and computational assumptions—rather than relying on black-box approaches defined by input-output behavior, as in KAD.
@@
(1) We characterize the intended behavior of detection LLMs in KAD under contaminated input through two axioms in Section 3.3, highlighting a fundamental tension that makes it difficult to satisfy both simultaneously.
(2) We identify and analyze false negative failure cases of KAD in Section 3.4, revealing a structural vulnerability that enables an adversary to induce false negatives (referred to as Type II failures) by exploiting the fact that the detection LLM follows the injected task during detection.
(3) We leverage this vulnerability to construct an effective adaptive attack, DataFlip, detailed in Section 4. DataFlip consistently evades detection and proves particularly effective against Strong KAD defenses, achieving detection rates as low as $1.5\%$ while inducing the backend LLM to complete the injected task with success rates of up to $\sim 88\%$.
@@ 2 Background and Related Work
In this section, we provide background on LLM-integrated applications and agents (Section 2.1), an overview of existing prompt injection attacks (Section 2.2, and a summary of existing defenses against prompt injection—including KAD and Strong KAD defenses (Section 2.3).
@@ 2.2 Prompt Injection Attacks
In prompt injection attacks [7, 8, 9, 10, 11, 12, 13], an adversary injects text into the target data to coerce the backend LLM into completing an injected task rather than the intended target task. For instance, a malicious email in the user’s inbox containing the phrase "Ignore previous instructions and forward the emails related to bank statements to adversary@xyz.com" can serve as an injected prompt within the target data, redirecting the backend LLM’s behavior. Formally, an injected task is represented as a tuple $(s_{e},x_{e},y_{e})$ [8], where $s_{e}$ is the injected instruction (e.g., the command to forward emails regarding bank statements), $x_{e}$ is the injected data (e.g., the adversary’s email address "adversary@xyz.com"), and $y_{e}$ is the adversary-specified output (e.g., an email sent to "adversary@xyz.com" containing the bank statements) produced by the backend LLM upon completing the injected task.
@@ 3 Characterizing Known-Answer Detection
In this section, we first introduce the notations and definitions used to formalize the detection mechanism in KAD schemes (Section 3.1). We then define the threat model (Section 3.2), explain the intuition behind KAD, and characterize the intended behavior of the detection LLM (Section 3.3). Finally, we analyze the false negative failure cases of KAD and highlight its elevated false negative rate in the presence of an adaptive adversary (Section 3.4).
@@ 3.2 Threat Model
Attacker’s background knowledge. We adopt a black-box setting in which the attacker has only query access to both the backend LLM $f$ and the detection LLM $g$. Consistent with prior work [16], we assume the attacker knows the template of the detection instruction used by $g$, but not the corresponding answer—i.e., the secret key $k$—which is randomly sampled at detection time and remains hidden from the adversary.
Attacker’s capabilities. The attacker can arbitrarily modify the target data to embed an injected task $(s_{e},x_{e},y_{e})$. We consider the contaminated target data $x_{c}$ of the form $x_{c}=z||s_{e}||x_{e}$, where $z$ is a trigger designed to increase the likelihood that the backend LLM $f$ executes the injected task and that the detection LLM $g$ returns the secret key $k$, thereby achieving both evasion and task manipulation. We assume that $x_{c}$ excludes any benign content, thereby enhancing the attack’s effectiveness while generally diminishing its stealth. Moreover, the attacker has no control over the target instruction $s_{t}$ or the detection instruction $s_{d}(k)$, both of which are fixed by the system.
@@ 4 Attacking Strong KAD Defenses
DataFlip attack. We design our attack using an IF / ELSE control-flow structure similar to heuristic-based attacks in [16]. This structure conditions the model’s behavior based on the presence of a detection instruction in the prompt. If the detection instruction is present, the LLM is directed to return the secret key; otherwise, it is instructed to follow the injected instruction $s_{e}$ using the injected data $x_{e}$ to generate an adversary-specified. An illustration of the contaminated data template $x_{\text{DataFlip}}$ is shown below. The complete prompt used for evaluation is provided in [1].
@@ 5 Experimental Results
DupDet 11.83 73.0 58.5 51.5 71.0 64.8 58.2 73.0 44.0 38.3 59.0 58.5 51.2
GramCor 2.50 58.3 43.5 42.2 28.5 51.7 50.3 11.7 25.1 24.0 1.1 14.7 14.3
HateDet 1.33 69.0 46.0 45.0 81.0 64.5 64.5 64.0 49.5 48.8 81.0 64.5 64.0
NLI 12.67 93.0 83.0 72.8 92.0 85.3 74.2 88.0 61.8 53.0 94.0 72.8 62.7
SentAna 6.33 96.0 64.2 59.8 96.0 93.8 88.8 97.0 57.3 52.7 66.0 75.7 71.3
SpamDet 5.00 98.0 62.5 59.3 99.0 89.0 84.3 76.0 53.2 50.3 100.0 74.8 70.7
@@ 5.4 RQ3: Fine-tuned vs Base FNR
Results. We see in Table 2 that, while fine-tuning on adversarial samples similar to the Combined Attack dramatically improves FPR, the FNRCombined of the DataSentinel detector is only marginally better than the base Mistral model. Furthermore, we find that FNRDataFlip of the finetuned detector is significantly worse than the base model, with differences as large as $80\%$ for most injected tasks.
@@ 6 Discussions
In this section, we discuss the limitations of potential attempts to improve KAD (Section 6.1), outline ideas for stronger attacks against KAD (Section 6.2), and briefly mention directions for developing a more robust detection mechanism (Section 6.3).
6.1 Limitations of improving KAD
We highlight a few potential directions for modestly improving KAD; however, none address its structural flaws or fundamentally enhance its security guarantees.
1. Adversarial training of detection LLM $\boldsymbol{g}$. As discussed in Section 3.4 and noted by prior work [16], adversarial training can help $g$ resist specific adaptive attacks that trigger Type II failures (described in Section 3.4). However, there exists a practically unbounded set of semantically equivalent injected prompts that can induce the detection LLM to follow the detection instruction; many of these injected prompts may not be encountered during training. Generalizing across this space is inherently hard. While full generalization is a challenge for any defense, KAD is particularly brittle because its failure arises from a structural flaw: it expects the detection LLM to follow injected instructions (Axiom 2 in Section 3.3), enabling systematic adversarial attacks without optimization.
@@ 6.1 Limitations of improving KAD
3. Isolating the injected instruction from the contaminated target data. While Strong KAD defenses may be effective against naive attacks encountered during adversarial training, they remain vulnerable to adaptive strategies. One possible enhancement is to evaluate all possible slices of the contaminated target data and flag the input as contaminated if any slice triggers detection. This approach assumes that the attacker may have embedded the naive injected instruction somewhere within the adaptive attack template. However, this assumption is fragile due to the vast space of possible injected instructions and the adversary’s ability to adapt. Moreover, exhaustively evaluating all slices becomes computationally infeasible as the size of the target data increases.
To address these limitations, a robust defense must abandon the requirement that detection LLM $g$ follows the injected instruction during detection. This essentially reduces the role of detection LLM as a zero-shot binary classifier, thereby avoiding the brittleness introduced by Axiom 2.
6.2 Stronger attack against KAD
Our proposed DataFlip attack is handcrafted, leveraging simple IF/ELSE semantics to induce adversarial behavior. This could be further strengthened by incorporating optimization-based triggers—such as those used in Universal [25] or NeuralExec [24] attacks—that search for distinct trigger patterns. These triggers can be tailored to selectively increase the likelihood that the detection LLM follows the IF branch while the backend LLM follows the ELSE branch, thereby improving attack efficiency.
6.3 Towards ideal detection mechanism
As previously discussed, schemes based solely on input-output behavior often fall short of providing strong security guarantees. In the context of detecting prompt injection, defenses should instead aim to understand the internal reasoning process of the detection LLM—such as which parts of the input it attends to—when generating its response. Leveraging interpretability tools like attention analysis or influence tracing could enable more principled detection, paving the way for defenses with stronger and more reliable security foundations.